OpenStudy (anonymous):

Law of Sines- The Ambiguous Case 1) m

OpenStudy (anonymous):



OpenStudy (anonymous):

I have to find m<A

OpenStudy (anonymous):

I don't get how to start. :/

OpenStudy (anonymous):

lets find the big triangle first. by the law of sines you have \[\frac{\sin(29)}{27}=\frac{\sin(B)}{31}\] \[\frac{31\sin(29)}{27}=\sin(B)\] \[\sin(B)=.5566\] rounded so \[B=\sin^{-1}(.5566)=33.82\]

OpenStudy (anonymous):